Find the slope of the line perpendicular to the graph of y=3/2x+3.
I don't have time to ask my teacher how to do it, and I missed a day so I wasn't there to get the notes and everything and when I went to the website I didn't understand the notes. So, I need some extra help here to help me figure out how to do this.

Answer by stanbon(75887)

Find the slope of the line perpendicular to the graph of y=3/2x+3.
------
y= (3/2)x+3 has a slope of 3/2
So any line perpendicular to it has a slope of (-3/2)
---------------------
Two slopes are perpendicular if m1*m2 = -1
=================================================
Cheers,
Stan H.